OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x + 3y = 1 3x - 4y = 10

hero (hero):

First solve both equations for y: y = (1-2x)/3 y = (10-3x)/-4 Second, set y = y: 3(10-3x) = -4(1-2x) 3. Distribute: 30 - 9x = -4 +8x Put like terms on one side: 17x = 34 Divide both sides by 17: x = 2